Skip to main contentSkip to solution

If E and F are two events such that P(E) = 0.8, P(F) = 0.7 and P(E ∩ F) = 0.6 then P(Eˉ/Fˉ)P(\bar{E}/\bar{F}) is:

Solution

✅ Correct Option: 3

P(E∪F)=0.8+0.7−0.6=0.9P(E\cup F)=0.8+0.7-0.6=0.9, so P(Eˉ∩Fˉ)=1−0.9=0.1P(\bar E\cap\bar F)=1-0.9=0.1. Also P(Fˉ)=0.3P(\bar F)=0.3. Hence P(Eˉ/Fˉ)=0.10.3=13P(\bar E/\bar F)=\frac{0.1}{0.3}=\frac{1}{3}.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question